what is second law of motion and how it is related to gravity


1
Expert's answer
2021-01-11T11:42:10-0500

Answer

According to second law of gravity "rate of change in momentum of body is directly proportion to the applied force on body."

So

"F=\frac{dP}{dt}\\and\\P=mv\\so\\F=\frac{d(mv)}{dt}\\m=constant\\then\\F=m\frac{dv}{dt}\\or\\F=ma\\"

If acceleration is due to gravity then

"a=g\\so\\F=mg"